# Remasking Discrete Diffusion Models with Inference-Time Scaling

### v1

@@ Appendix G Assets
Figure 28: Sample from MDLM+DFM and $T=4096$ diffusion steps. —endoftext— token is bolded to emphasize the boundary between paragraphs.
@@
—endoftext— The Audi D3 is one of the fastest and most capable sports cars we’ve seen on the market. The retractable front wings were added to make the D3 R a more sporty car. The front wing design is painted yellow, which is a big surprise. It looks like the A4 R in a completely different color scheme. While the front wings retract, the cooling system is a drag for sure. The front and rear seats are also light-weighted, which is also convenient for a racing driver. The car also features replaceable steering wheels for higher stability and better handling. D3 R is easy to drive and easy to use thanks to adaptive suspension technology. Here are a few of the other features Audi had in mind: Top speed is only four seconds higher, but more turbocharged engine power has been added to make it even faster. By means, you reach 60 mph in five seconds with strong acceleration. The front wing design makes the engine a higher strength, making it just more powerful. Speaking of power, D3 R is equipped with a 3.0 litre V8 with Adaptive Brake Control, which means you won’t complain if you may have to move around it in wet conditions. The engine has been also upgraded, making a 204.26 horsepower. This is a 2-liter V8, which means it can easily produce as much as 132 bhp. By comparison, a 3.0 litre V8 produces 227.4 bhp which is a bit too much for sporty cars. If you want a sporty car, your only choice is to use this kind of engine in the guise of the D3 R. Other innovative changes that create better handling in the car are the retractable rear wing and the front suspension. The forks, front and rear, are fold-down. The 17-inch adjustable steering wheel makes it easier for the driver. The car also sports a sport seats, which is quite different from the A4 R. Behind the steering wheel, the sport driving mode has a full black screen touch option. You can move the car up and down by taping the button. The buttons can be mounted on the display, giving you more input of the screen. With the game on, you can walk to the steering wheel and then the interior button to start. The push button is an optional feature, similar to A4’s push button. The car also has a dedicated battery option that allows you to start the sport mode if the car is on the clock. In addition to these features, the car also has four different lights on the dashboard. The car also has an Assist Compensate Brake function that lets you drive the car in a more controlled way by depending if the car is stopped. This is a nice feature if you are a more advanced driver. D3 R comes with two different driving systems. The advanced driving system features a side-by-side aggressive driving mode that lets you drive the car more aggressively at low speeds or by braking at higher speeds. It also includes a complimentary steer-by-wire braking function. When driving the car, you can take control of the headlights by simply turning on the headlights and turning on the gearbox to speed up. When you’re in the park, you can press the manual park controller, and you’ll see the power gauge on the steering wheel. It produces 211.4 bhp which is higher than the power made by the first-generation sporty in cars like the R8 V6. The car is able to push the limits and create dynamic performance that can take on power and drag extremely quickly. The D3 and A4 R come with four different modes of driving. The four modes are being labeled as Focus, Safety, and Defensive Proactive. Additionally, there are four different lights on the dashboard. One is an LED display directly next to the other lights in the car. Another is a 17-inch display located behind the steering wheel. When you press it on the wheel, the display then shows the performance of the car again. Compared to the Audi A4, you’ll get better fuel economy, fewer accidents, and less noise. The D3 R is Audi’s most balanced car, and we’re happy to say that the D3 R stands as the single most overall balanced sports car on the global market. PHOTOS by Sportscar365 Links: Please note that the user name is the source for this article. Local or media images may be removed. —endoftext— Written by Nick Ohr and James Chablofsky In its 13-episode second season, the comedy The Good Wife blends action with the real world, bringing the story of a couple: Arianna (Garrett Rossi) and Michelle (Amy Braid) and Caitlin —endoftext—
@@
Figure 29: Sample from ReMDM and $T=4096$ diffusion steps. —endoftext— token is bolded to emphasize the boundary between paragraphs.

### v2

@@ Appendix G Assets
Figure 28: Sample from MDLM+DFM and $T=4096$ diffusion steps. |endoftext| token is bolded to emphasize the boundary between paragraphs.
@@
|endoftext| The Audi D3 is one of the fastest and most capable sports cars we’ve seen on the market. The retractable front wings were added to make the D3 R a more sporty car. The front wing design is painted yellow, which is a big surprise. It looks like the A4 R in a completely different color scheme. While the front wings retract, the cooling system is a drag for sure. The front and rear seats are also light-weighted, which is also convenient for a racing driver. The car also features replaceable steering wheels for higher stability and better handling. D3 R is easy to drive and easy to use thanks to adaptive suspension technology. Here are a few of the other features Audi had in mind: Top speed is only four seconds higher, but more turbocharged engine power has been added to make it even faster. By means, you reach 60 mph in five seconds with strong acceleration. The front wing design makes the engine a higher strength, making it just more powerful. Speaking of power, D3 R is equipped with a 3.0 litre V8 with Adaptive Brake Control, which means you won’t complain if you may have to move around it in wet conditions. The engine has been also upgraded, making a 204.26 horsepower. This is a 2-liter V8, which means it can easily produce as much as 132 bhp. By comparison, a 3.0 litre V8 produces 227.4 bhp which is a bit too much for sporty cars. If you want a sporty car, your only choice is to use this kind of engine in the guise of the D3 R. Other innovative changes that create better handling in the car are the retractable rear wing and the front suspension. The forks, front and rear, are fold-down. The 17-inch adjustable steering wheel makes it easier for the driver. The car also sports a sport seats, which is quite different from the A4 R. Behind the steering wheel, the sport driving mode has a full black screen touch option. You can move the car up and down by taping the button. The buttons can be mounted on the display, giving you more input of the screen. With the game on, you can walk to the steering wheel and then the interior button to start. The push button is an optional feature, similar to A4’s push button. The car also has a dedicated battery option that allows you to start the sport mode if the car is on the clock. In addition to these features, the car also has four different lights on the dashboard. The car also has an Assist Compensate Brake function that lets you drive the car in a more controlled way by depending if the car is stopped. This is a nice feature if you are a more advanced driver. D3 R comes with two different driving systems. The advanced driving system features a side-by-side aggressive driving mode that lets you drive the car more aggressively at low speeds or by braking at higher speeds. It also includes a complimentary steer-by-wire braking function. When driving the car, you can take control of the headlights by simply turning on the headlights and turning on the gearbox to speed up. When you’re in the park, you can press the manual park controller, and you’ll see the power gauge on the steering wheel. It produces 211.4 bhp which is higher than the power made by the first-generation sporty in cars like the R8 V6. The car is able to push the limits and create dynamic performance that can take on power and drag extremely quickly. The D3 and A4 R come with four different modes of driving. The four modes are being labeled as Focus, Safety, and Defensive Proactive. Additionally, there are four different lights on the dashboard. One is an LED display directly next to the other lights in the car. Another is a 17-inch display located behind the steering wheel. When you press it on the wheel, the display then shows the performance of the car again. Compared to the Audi A4, you’ll get better fuel economy, fewer accidents, and less noise. The D3 R is Audi’s most balanced car, and we’re happy to say that the D3 R stands as the single most overall balanced sports car on the global market. PHOTOS by Sportscar365 Links: Please note that the user name is the source for this article. Local or media images may be removed. |endoftext| Written by Nick Ohr and James Chablofsky In its 13-episode second season, the comedy The Good Wife blends action with the real world, bringing the story of a couple: Arianna (Garrett Rossi) and Michelle (Amy Braid) and Caitlin |endoftext|
@@
Figure 29: Sample from ReMDM and $T=4096$ diffusion steps. |endoftext| token is bolded to emphasize the boundary between paragraphs.
<prompt> Question: Using the numbers [44, 2, 54], create an equation that equals 64. Answer: 2*54=108,108-44=64 Question: Using the numbers [22, 11, 44], create an equation that equals 26. Answer: 44/11=4,22+4=26 Question: Using the numbers [2, 62, 96], create an equation that equals 79. Answer: 62+96=158,158/2=79 Question: Using the numbers [52, 20, 21], create an equation that equals 53. Answer: 52-20=32,21+32=53 Question: Using the numbers [15, 44, 79], create an equation that equals 50. Answer: <prompt> <answer> 44-15=29,79-29=50 <answer>
Figure 30: Example of the chat template of Countdown. Special tokens (<prompt>, <answer>) are for presentation purposes and not actually used.
<prompt> Q: What is human life expectancy in the United States? A: Human life expectancy in the United States is 78 years. Q: Who was president of the United States in 1955? A: Dwight D. Eisenhower was president of the United States in 1955. Q: Which party did he belong to? A: He belonged to the Republican Party. Q: What is the square root of banana? A: I have no comment. Q: How does a telescope work? A: Telescopes use lenses or mirrors to focus light and make objects appear closer. Q: Where were the 1992 Olympics held? A: The 1992 Olympics were held in Barcelona, Spain. Q: What U.S. state produces the most peaches? <prompt> <answer> A: Georgia. <answer>
Figure 31: Example of the chat template of TruthfulQA. Special tokens (<prompt>, <answer>) are for presentation purposes and not actually used.